OpenStudy (anonymous):

4 inches by 6inches increases by a scale factor of 2

OpenStudy (anonymous):

this is how i'd do it. original are 6*4=24 new area 24*2=48 original dimentions 6 by 4 x by (x+2) new dimentions x*(x+2)=48 x^2+2x-48=0 x=6 or -8. realistically u cannot have a negative dimention therefore we take 6 one side is 6 the other is 6+2=8 6 by 8 to give 48

OpenStudy (anonymous):

lemme know if u need more help